# THeGCN: Temporal Heterophilic Graph Convolutional Network

### v1

@@ Introduction
Appendix
Construction of Datasets
@@
For $3$ traffic datasets, the raw data is collected from the Caltrans Performance Measurement System (PeMS) (Chen et al. 2001), which is composed of the freeway system in major areas of California. The traffic statistics from three areas are chosen: Bay Area (Station 4), Los Angeles (Station 7), and San Diego (Station 11) to build the corresponding PEMSBA, PEMSLA, and PEMSSD datasets in our paper. In detail, for each dataset, each sensor is considered as a node. A 120-minute interval traffic flow data from each sensor ($v_{i}$ at time $t$) is collected. The total number of intervals is 84 (i.e., one week). Each interval contains three aspects of information: flow rate, speed, and occupancy of the sensor. We also have the location (i.e., latitude and longitude) of each sensor. For each $v_{i}$ at time $t$, we first select 5 nodes with the most similar flow rate. Then, a threshold is set (i.e., $\tau=0.000001$) to filter out sensors/nodes with Euclidean distance to $v_{i}$ larger than the threshold. The remaining sensors $v_{j}$s within the distance threshold form the interactions with $v_{i}$ at $t$. After constructing the interactions, we divide the range of speed values into 10 intervals with equal length. Depending on the interval of the speed value of $v_{i}$ at $t$, a 10 dimension one hot vector is built as the attribute $\mathbf{x}^{t}_{v_{i}}$, where only one entry is 1 and the remaining entries are 0. At last, we utilize the occupancy to build the labels of nodes. For all nodes at different moments, we collect the occupancy values of them during the last 12 intervals (i.e., the last day). Similar to the process of constructing node attributes. The occupancy values are divided into 5 intervals with the same length and each interval corresponds to one class. For these labeled nodes, the edge/interaction heterophily ranges from 34.61% of PEMSLA to 68.04% of PEMSSD. In addition, a metric named temporal changing ratio is proposed. This metric measures the ratio of nodes whose labels have changed during the last 12 intervals. We can see that over 95% of nodes have changed the labels during this period, which also matches the motivation of this work.
@@
The Reddit dataset is generated from Reddit. The raw data of Reddit contains web content ratings, discussions among users, and some social news. To construct the Reddit dataset in our experiments, The graph is constructed by viewing posts as nodes. If two posts at the same time intervals share the same keywords, we build an interaction between them. We apply word2vec (Mikolov et al. 2013) on the post comments to generate the node attributes. The labels of nodes are constructed by (Fan, Yao, and Joe-Wong 2021) and remain the same during the whole time period. The Brain dataset is generated from functional magnetic resonance imaging (fMRI) data  [^9] https://tinyurl.com/y4hhw8ro. Nodes represent cubes of brain tissue, and two nodes are connected during a time period if they show similar degrees of activation during that time period. We randomly split labeled nodes  [^10] In PEMSBA/PEMSLA/PEMSSD, since labels for the same node change with time, the same node with different labels at different times are viewed as two labeled nodes. In Reddit, the same node only appears once in the training/validation/test set. in every dataset into 60/20/20% for training, validation, and testing.

### v2

@@ 6 Conclusion and Limitation
7 Appendix
7.1 Construction of Datasets
@@
For $3$ traffic datasets, the raw data is collected from the Caltrans Performance Measurement System (PeMS) Chen et al. (2001), which is composed of the freeway system in major areas of California. The traffic statistics from three areas are chosen: Bay Area (Station 4), Los Angeles (Station 7), and San Diego (Station 11) to build the corresponding PEMSBA, PEMSLA, and PEMSSD datasets in our paper. In detail, for each dataset, each sensor is considered as a node. A 120-minute interval traffic flow data from each sensor ($v_{i}$ at time $t$) is collected. The total number of intervals is 84 (i.e., one week). Each interval contains three aspects of information: flow rate, speed, and occupancy of the sensor. We also have the location (i.e., latitude and longitude) of each sensor. For each $v_{i}$ at time $t$, we first select 5 nodes with the most similar flow rate. Then, a threshold is set (i.e., $\tau=0.000001$) to filter out sensors/nodes with Euclidean distance to $v_{i}$ larger than the threshold. The remaining sensors $v_{j}$s within the distance threshold form the interactions with $v_{i}$ at $t$. After constructing the interactions, we divide the range of speed values into 10 intervals with equal length. Depending on the interval of the speed value of $v_{i}$ at $t$, a 10 dimension one hot vector is built as the attribute $\mathbf{x}^{t}_{v_{i}}$, where only one entry is 1 and the remaining entries are 0. At last, we utilize the occupancy to build the labels of nodes. For all nodes at different moments, we collect the occupancy values of them during the last 12 intervals (i.e., the last day). Similar to the process of constructing node attributes. The occupancy values are divided into 5 intervals with the same length and each interval corresponds to one class. For these labeled nodes, the edge/interaction heterophily ranges from 34.61% of PEMSLA to 68.04% of PEMSSD. In addition, a metric named temporal changing ratio is proposed. This metric measures the ratio of nodes whose labels have changed during the last 12 intervals. We can see that over 95% of nodes have changed the labels during this period, which also matches the motivation of this work.
@@
The Reddit dataset is generated from Reddit. The raw data of Reddit contains web content ratings, discussions among users, and some social news. To construct the Reddit dataset in our experiments, The graph is constructed by viewing posts as nodes. If two posts at the same time intervals share the same keywords, we build an interaction between them. We apply word2vec Mikolov et al. (2013) on the post comments to generate the node attributes. The labels of nodes are constructed by Fan et al. (2021) and remain the same during the whole time period. The Brain dataset is generated from functional magnetic resonance imaging (fMRI) data  [^9] https://tinyurl.com/y4hhw8ro. Nodes represent cubes of brain tissue, and two nodes are connected during a time period if they show similar degrees of activation during that time period. We randomly split labeled nodes  [^10] In PEMSBA/PEMSLA/PEMSSD, since labels for the same node change with time, the same node with different labels at different times are viewed as two labeled nodes. In Reddit, the same node only appears once in the training/validation/test set. in every dataset into 60/20/20% for training, validation, and testing.